cooling: 2) inner heat screen - C-side: leak test, A-side hoses+leak test 4) leak and pressure test (up to 300 mbar) for all sectors: 70% done, rest today 5) NPT adapters for pressure relief valves: 14.12. delivery. Pressure gauges end Jan 6) ss adapters for all cooling connections: offer not yet received 6) leak hunting: 2x chamber cooling circuits still not ok (A, C) FEC: 1) new grounding scheme: change cables 3) Wiener test: this week cabling: 1) C12 gating test: open (Danilo) 2) connect patch panel cables (cover, edge) Rainer: redo 1 cable 3) redo skirt cable routing and testing: open, part in cavern 07 4) storage and fixation of cable rolls: Rainer, technicians: ongoing ... 5) CAN-ELMB cable: replace with shielded: Rainer --> 07 6) ELMB adapter resistor: Marco: do connector, test 7) connector at bretelle display: Thomas: next week misc: 2) ITS cameras: 18.12., Bjorn 3) orientation of bretelles (Spanner)? 4) remove HV cable: Joachim